# Mechanistic Fine-tuning for In-context Learning

### v1

@@ A.3 Details of PID Algorithm
On each model update step (i.e., when the gradients from all the samples of the -th pseudo batch (of data samples) are propagated), we calculated the identified induction head numbers from the induction head filter described in §3 and Fig. 1 averaged on the data samples as . Given the similar averaged induction head number on the previous time step as , we can calculate the updated term555Remind that for the sake of simplicity, we only control , given the findings of and are antagonistic, as shown in §5.3. by standard PID algorithm as:
@@ A.4 Experiment Protocol of Unseen Label
In Fig. LABEL:fig:unseen, we examine that ABFT model can utilize the demonstration with wrong label to improve the ICL performance. Here we introduce the experiment protocol.
@@ Appendix C Augmentation Results
C.1 Augmentation Results for Data Efficiency (Fig. LABEL:fig:data_eff)
We repeat the data efficiency experiments shown in Fig. LABEL:fig:data_eff on Qwen2.5 32B, as shown in Fig. 15. The results are globally consistent with Fig. LABEL:fig:data_eff.

### v2

@@ A.3 Details of PID Algorithm
On each model update step (i.e., when the gradients from all the samples of the -th pseudo batch (of data samples) are propagated), we calculated the identified induction head numbers from the induction head filter described in §3 and Fig. 1 averaged on the data samples as . Given the similar averaged induction head number on the previous time step as , we can calculate the updated term666Remind that for the sake of simplicity, we only control , given the findings of and are antagonistic, as shown in §5.3. by standard PID algorithm as:
@@ A.4 Experiment Protocol of Unseen Label
In Fig. 3, we examine that ABFT model can utilize the demonstration with wrong label to improve the ICL performance. Here we introduce the experiment protocol.
@@ Appendix C Augmentation Results
C.1 Augmentation Results for Data Efficiency (Fig. 3)